OpenStudy (anonymous):

Find the highest point on the cardioid r=1+cos(theta)

OpenStudy (aum):

Need to maximize the y-component or \(r * \sin(\theta) = \sin(\theta)(1 + \cos(\theta))\)

OpenStudy (anonymous):

now where does rsin(theta) come from?

OpenStudy (aum):

Polar coordinates to Cartesian coordinates. x = r * cos(theta) y = r * sin(theta)

OpenStudy (aum):

Highest point is when y is maximum. \(y = r * \sin(\theta) = \sin(\theta)(1 + \cos(\theta))\) Set \(\large \frac{dy}{d\theta} = 0\) and solve for \(\theta\). Find which \(\theta\) would give the highest y. Put that \(\theta\) in r and calculate r.